Cobb Douglas
U=X1/2 1,X1/2 2) budget constraints of M=P1x1+p2x2 determine the marginal utilities of X1 x2 determine the quantity demanded for good 1 + 2
1
Expert's answer
2021-08-02T14:58:52-0400

Cobb Douglass Utility Function: "U=X_1^{\frac{1}{2}}X_2^{\frac{1}{2}}"

Budget Constraint: "M=P_1X_1+P_2X_2"

Marginal Utility is the utility derived by an additional consumption of a unit of a good.

 

Marginal Utility of X1:

"MU_{X_1}=\frac{\delta U}{\delta X_1}\\

MU_{X_1}=\frac{dU}{dX_1}\\

U=X_1^{\frac{1}{2}}X_2^{\frac{1}{2}}\\

\frac{dU}{dX_1}=\frac{1}{2}X_1^{\frac{-1}{2}}X_2^{\frac{1}{2}}\\

\frac{dU}{dX_1}=\frac{1}{2}(\frac{X_2}{X_1})^{\frac{1}{2}}"

So, the Marginal Utility of X1 is "\frac{1}{2}(\frac{X_2}{X_1})^{\frac{1}{2}}"

 

Similarly, we can find the marginal utility of X2:

"MU_{X2}=\frac{\delta U}{\delta X2}\\

MU_{X2}=\frac{dU}{dX2}\\

U=X1^{\frac{1}{2}}X2^{\frac{1}{2}}\\

\frac{dU}{dX2}=\frac{1}{2}X1^{\frac{1}{2}}X2^{\frac{-1}{2}}\\

\frac{dU}{dX2}=\frac{1}{2}(\frac{X1}{X2})^{\frac{1}{2}}"

So, Marginal Utility of X1 is "\frac{1}{2}(\frac{X1}{X2})^{\frac{1}{2}}"


For quantity demand, we need the following condition:

"\frac{MU_{X_1}}{MU_{X_2}}=\frac{P_1}{P_2}\\

\frac{\frac{1}{2}(\frac{X_2}{X_1})^{\frac{1}{2}}}{ \frac{1}{2}(\frac{X_1}{X_2})^\frac{1}{2}}= \frac{P_1}{P_2}"

Simplifying the above expression, we get

"\frac{X_2}{X_1}=\frac{P_1}{P_2}\\

P_1X_1=P_2X_2"

Putting this in the budget constraint

Budget Constraint: "M=P_1X_1+P_2X_2"

"P_1X_1+P_2X_2=M\\

Since\space P_1X_1=P_2X_2\\

P_1X_1+P_1X_1=M\\

2P_1X_1=M\\

X_1=\frac{M}{2P_1}"

So, the demand for X1 is"X_1=\frac{M}{2P_1}"

 

We will substitute the value of X1 in "P_1X_1=P_2X_2" to find out the demand for X2

"P_1X_1=P_2X_2\\

X_2=(\frac{P_1}{P_2})\times X_1"

Putting the value of "X_1=\frac{M}{2P_1}"

"X_2=(\frac{P_1}{P_2})\times(\frac{M}{2P_1})\\

X_2=\frac{M}{2P_2}"

So, the demand for X2 is "X_2=\frac{M}{2P_2}"

Quantity demanded of good "1+2: X_1+X_2"

"X_1+X_2=(\frac{M}{2P_1})+(\frac{M}{2P_2})"